# Bipartite Fluctuations of Critical Fermi Surfaces

### v2

@@ Abstract
Fluctuations of conserved quantities within a subsystem are non-local observables that provide unique insights into quantum many-body systems. In this paper, we explore the behaviors of bipartite charge (and spin) fluctuations across interaction-driven “metal-insulator transitions” out of Landau Fermi liquids. The “charge insulators” include a class of non-Fermi-liquid states of fractionalized degrees of freedom, such as compressible composite Fermi liquids (for spinless electrons) and incompressible spin-liquid Mott insulators (for spin-$1/2$ electrons). We find that charge fluctuations $\mathcal{F}$ exhibit distinct leading-order scalings across the phase transition: $\mathcal{F}\sim L\log(L)$ in Landau Fermi liquids and $\mathcal{F}\sim L$ in charge insulators, where $L$ is the linear size of the subsystem under bipartition. In the case of composite Fermi liquids (under certain conditions), we also identify a universal constant term $-\textrm{f}(\theta)|\sigma_{xy}|/(2\pi)$ arising when the subsystem geometry incorporates a sharp corner. Here, $\textrm{f}(\theta)$ represents a function of the corner angle, and $\sigma_{xy}$ denotes the Hall conductivity. At the critical point of each example, provided the transition is continuous, we find that the leading-order scaling $\mathcal{F}\sim L$ is accompanied by a subleading universal corner contribution $-\log(L)\textrm{f}(\theta)C_{\rho}/2$ with the same angle dependence $\textrm{f}(\theta)$. The universal coefficient $C_{\rho}$ is linked to the predicted universal longitudinal (and Hall) resistivity jump $\Delta\rho_{xx}$ (and $\Delta\rho_{xy}$) at the critical point.
@@ I Introduction
Measuring bipartite fluctuations of local observables provides a powerful approach to characterizing entanglement properties, symmetries, and correlations in various quantum phases and phase transitions. Although entanglement has become a fundamental organizing principle for the study of quantum matter [1, 2], experimental measurement of many-body entanglement remains a challenge. Significant progress (see, e.g., Ref. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]) has been made toward establishing a direct relationship between entanglement entropy and the fluctuations of globally conserved quantities, such as particle number and magnetization. This relationship has been demonstrated to be feasible in special cases, such as one-dimensional Luttinger liquids [6, 8] and free fermions in higher dimensions [9, 8, 4]. However, the connection between the two in strongly correlated systems beyond these examples is still poorly understood. The two quantities are observed to differ from each other in continuous symmetry-breaking phases, where the entanglement entropy still follows a boundary-law scaling while the bipartite fluctuations exhibit a multiplicative logarithmic enhancement to the boundary law. Building on this observation, Ref. [13] proposed the use of the distinct scalings of bipartite fluctuations to detect conventional Landau phase transitions.
@@
Despite decades of study, understanding unconventional quantum phases and phase transitions that are beyond Landau’s symmetry-breaking paradigm remains a central problem in condensed matter physics. In recent years, conceptual breakthroughs [14, 15] have been achieved by employing extended operators to define generalized symmetries and to characterize phases and phase transitions (see Ref. [16, 17] and references therein). For instance, abelian topological orders can be understood in terms of the condensation of 1-dimensional objects, leading to the spontaneous breaking of discrete $1$-form symmetries. Similarly, the Coulomb phase of electrodynamics can be interpreted as a spontaneous symmetry-breaking phase of $\textrm{U}(1)$ $1$-form symmetry, where gapless photons serve as Goldstone modes. This helps establish a generalized Landau symmetry paradigm. In this case, the criterion for determining whether the phase is symmetric or not is given by the scaling of disorder operators (or Wilson loops under duality). In this context, the concept of bipartite fluctuations is again closely related, which can be identified as the $\textrm{U}(1)$ disorder operator under the small-angle limit [18, 19].
@@
While extended operators play an important role in the conceptual understanding of phases of matter, recent studies have also shed light on intriguing quantitative aspects of disorder operators (and related bipartite fluctuations), particularly their scaling behaviors at conformally invariant quantum critical points [20, 18, 19, 21, 22, 23, 24, 25, 26]. For instance, when considering the configuration of the subsystem as depicted in Fig. 1, bipartite fluctuations exhibit a universal corner contribution with logarithmic scaling. This contribution has a universal angle dependence and is directly proportional to the current central charge $C_{J}$ of the conformal field theory (CFT) [20, 18, 19, 21]. Given the sometimes uncertain fate of proposed lattice realizations for exotic quantum critical points, numerical simulations, such as quantum Monte Carlo simulations, can assist in identifying whether these lattice models correspond to unitary CFTs by examining the sign of the universal corner contribution [22, 23, 24, 25, 26]. Moreover, Ref. [21] showed that the corner term of bipartite fluctuations in isotropic integer quantum Hall (QH) and Laughlin states is determined by the DC Hall conductivity.
@@
Another notable example that has appeared in literature is the bandwidth-tuned transition from a Landau Fermi liquid (FL) to a composite Fermi liquid (CFL) describing the half-filled Landau level. The theoretical possibility of a continuous transition was originally proposed by Ref. [55] and later refined by Ref. [56] targeting TMD Moiré materials [^2] The existence of CFLs at 1/2 and 1/4 fillings in twisted $\textrm{MoTe}_{2}$ is supported by Ref. [57, 58].. A general criterion [59, 60, 61, 62] distinguishing charge insulators from metals is the presence of a finite localization tensor, i.e., finite polarization fluctuations. Accordingly, as will be illustrated below, the CFL may be viewed as a “charge insulator,” and the CFL-FL transition can be interpreted as a “metal-insulator transition” in a broad sense.
All these “metal-insulator transitions” share a common theoretical framework, which we will review in Sec. III, where electrons are fractionalized into fermionic and bosonic partons. The fermion sector forms a stable Fermi-surface state, while the boson sector contains a single relevant operator that drives the transition from a superfluid to a gapped state. Crucial insight into the possible transitions comes from the filling constraints of the bosons under translation symmetry and the Lieb-Schultz-Mattis theorem [54, 56]. Without better terminology, we refer to this type of quantum critical points as critical Fermi surfaces, following Refs. [50, 51].
Figure 1: The real-space subsystem $\mathcal{A}$ with only one corner angle $0<\theta<\pi$. The linear size of $\mathcal{A}$ is denoted by $L$.
@@
In contrast to conformally invariant quantum critical points, many powerful theoretical tools, such as the conformal bootstrap [64], are no longer applicable to strongly correlated metals. However, it is theoretically feasible in Monte Carlo simulations [22, 23, 24, 25, 26] to extract the subleading universal corner term of $\mathcal{F}$ using the method proposed in Ref. [65]. We anticipate that our findings, which establish a connection between the universal data $C_{\rho}$ of the critical points and charge fluctuations which are numerically and experimentally accessible, could aid future studies in identifying the existence and lattice realizations of these exotic quantum critical points in metals.
@@ II Bipartite Fluctuations
In this section, we begin with a brief introduction to the preliminaries of bipartite fluctuations in various quantum states of matter, along with a general criterion for distinguishing charge insulators from metals. We then summarize our main findings on bipartite fluctuations across a class of bandwidth-tuned continuous “metal-insulator transitions.”
@@ II.1 Preliminaries I
For any many-body system with a $\textrm{U}(1)$ global symmetry, the bipartite fluctuations can be defined as
$\mathcal{F}_{\mathcal{A}}=\left\langle\left(\int_{\mathcal{A}}J^{\tau}\right)-\left\langle\int_{\mathcal{A}}J^{\tau}\right\rangle\right\rangle^{2}.$ (1)
Here, $J^{\tau}$ represents the zeroth component of the conserved Noether current $J^{\mu}$ satisfying $\partial_{\mu}J^{\mu}=0$, and $\mathcal{A}$ is a spatial subregion within the total system. The expectation value is taken with respect to the ground state at zero temperature. By definition, the large-scale scaling behavior of $\mathcal{F}_{\mathcal{A}}$ is ultimately governed by the long-wavelength behavior of the equal-time density-density correlation $\Pi^{\tau\tau}=\langle J^{\tau}J^{\tau}\rangle$. In App. D, we undertake the task of evaluating $\mathcal{F}_{\mathcal{A}}$ assuming a generic power-law instantaneous charge correlation
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})=\frac{C_{0}}{|\bm{x}|^{\alpha}}$ (2)
in two dimensions. The results are summarized in Tab. 1.
@@
$32<\alpha<$ 3 $L^{4-\alpha}$ $L^{4-\alpha}$
@@
$43<\alpha<$ 4 $L/\epsilon^{\alpha-3}$ $L^{4-\alpha}$
@@
$\alpha>4$ $L/\epsilon^{\alpha-3}$ 00
Table 1: The scaling behaviors of the bipartite fluctuations $\mathcal{F}_{\mathcal{A}}$ based on the density-density correlation Eq. (2). Here, $L$ represents the linear size of $\mathcal{A}$, and $\epsilon$ is a gauge-invariant UV cut-off introduced in the vortex theory (see App. D). There are some notable features. The leading-order term exhibits boundary-law scaling when $\alpha>3$. A universal subleading logarithmic term appears only when $\alpha=4$ and the geometry of $\mathcal{A}$ includes any sharp corners. For $\alpha>4$, the universal subleading term vanishes in the large-$L$ limit.
The simplest compressible state is given by the free fermi gas. Utilizing the free-fermion propagator, one can easily obtain $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})$ in real space or the static structure factor $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ in momentum space
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})$ $=\frac{k_{F}}{4\pi^{3}}\frac{1}{|\bm{x}|^{3}},$
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ $=-\frac{k_{F}|\bm{k}|}{2\pi^{2}},$ (3)
where $k_{F}$ denotes the fermi momentum. According to Tab. 1, we observe the scaling $\mathcal{F}\sim L\log(L)$. This is a well-known result and is closely related to entanglement entropy [66, 67, 68, 69, 9, 8, 4]. In App. A, we reproduce Eq. (3) based the $\textrm{LU}(1)$ anomaly of Fermi surface states [45], which holds even in the presence of strong forward scattering preserving the $\textrm{LU}(1)$ symmetry. The bonus is a geometric interpretation [70] of the static structure factor through the area difference when shifting the Fermi surface, as depicted in Fig. 2.
Figure 2: The geometric interpretation of the static structure factor $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{q})$ for Fermi surface states with LU(1) symmetry, where the shaded area represents its value.
Another simple compressible state is provided by the superfluid phase. Let us denote the $\textrm{U}(1)$ order parameter by $\phi$. The gapless Goldstone mode $\theta$ corresponds to the angle fluctuation of $\phi=\rho_{s}e^{\mathtt{i}\theta}$, where $\rho_{s}$ represents the stiffness. Since the power counting of the action $\int\textrm{d}^{3}x(\partial\theta)^{2}$ yields a scaling dimension of $\Delta[\partial\theta]=3/2$ for the $\textrm{U}(1)$ charge density, the corresponding equal-time correlation scales as $|\bm{x}|^{-3}$. Once again, one finds the scaling of bipartite fluctuations $\mathcal{F}\sim L\log(L)$.
@@
=#$\mathcal{F}=\#L-\frac{C_{J}}{2}\textrm{f}(\theta)\log(L)+\ldots.$ (5)
where ##\## is a non-universal number depending on the UV cut-off, and the universal angle dependence
@@
applies to any $(2+1)$-dimensional CFTs. This corner contribution has also been previously discussed in the context of non-interacting Dirac systems [20]. However, the question remains whether the universal behavior described in Eq. (5) could also exist at quantum critical points lacking conformal symmetry.
Insulators with a charge gap represent another standard type of incompressible states. Due to the exponential decay of the equal-time density-density correlation, the leading-order scaling obeys the boundary law. An intriguing observation from Ref. [21] is that certain gapped QH insulators, which possess continuous translational and rotational symmetries, exhibit a universal corner contribution when considering the subregion geometry shown in Fig. 1:
=#$\mathcal{F}=\#L-\frac{|\sigma_{xy}|}{2\pi}\textrm{f}(\theta)+\ldots.$ (7)
The angle dependence is once again given by Eq. (6), and $\sigma_{xy}=\nu/(2\pi)$ represents the DC Hall conductivity, where $\nu$ denotes the filling factor. The corner term for integer fillings $\nu\in\mathbb{Z}$ can be derived analytically, while the fractional QH insulator at $\nu=1/3$ has been verified using Monte Carlo simulations based on the Laughlin wavefunction [21]. It would be interesting to explore whether a similar corner term exists in compressible QH states at even-denominator fillings and to understand the conditions under which the formula in Eq. (7) holds.

### v3

@@ Abstract
Fluctuations of conserved quantities within a subsystem are non-local observables that provide unique insights into quantum many-body systems. In this paper, we study bipartite charge (and spin) fluctuations across interaction-driven “metal-insulator transitions” out of Landau Fermi liquids. The “charge insulators” include a class of non-Fermi-liquid states of fractionalized degrees of freedom, such as compressible composite Fermi liquids (for spinless electrons) and incompressible spin-liquid Mott insulators (for spin-$1/2$ electrons). We find that charge fluctuations $\mathcal{F}$ exhibit distinct leading-order scalings across the transition: $\mathcal{F}\sim L\log(L)$ in Landau Fermi liquids and $\mathcal{F}\sim L$ in charge insulators, where $L$ is the linear size of the subsystem. In composite Fermi liquids, under certain conditions, we also identify a universal constant term $\textrm{f}(\theta)|\sigma_{xy}|/(2\pi)$ when the subsystem geometry contains a sharp corner, where $\textrm{f}(\theta)$ denotes a function of the corner angle, and $\sigma_{xy}$ is the Hall conductivity. At the critical point, provided the transition is continuous, the leading scaling $\mathcal{F}\sim L$ is accompanied by a subleading universal corner contribution $\log(L)\textrm{f}(\theta)C_{\rho}/2$ with the same angle dependence $\textrm{f}(\theta)$, and the universal coefficient $C_{\rho}$ is directly related to the predicted universal jumps in longitudinal and Hall resistivities. These results establish fluctuation-transport relations, paving the way for numerical and experimental studies of unconventional quantum criticalities in metals.
@@ I Introduction
Measuring bipartite fluctuations of local observables provides a powerful approach to characterizing entanglement properties, symmetries, and correlations in various quantum phases and phase transitions. Although entanglement has become a fundamental organizing principle for the study of quantum matter [1, 2], experimental measurement of many-body entanglement remains a challenge. Significant progress (see, e.g., Refs. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]) has been made toward establishing a direct relationship between entanglement entropy and the fluctuations of globally conserved quantities, such as particle number and magnetization. This relationship has been demonstrated to be feasible in special cases, such as one-dimensional Luttinger liquids [6, 8] and free fermions in higher dimensions [9, 8, 4]. However, the connection between the two in strongly correlated systems beyond these examples is still poorly understood. The two quantities are observed to differ from each other in continuous symmetry-breaking phases, where the entanglement entropy still follows a boundary-law scaling while the bipartite fluctuations exhibit a multiplicative logarithmic enhancement to the boundary law. Building on this observation, Ref. [13] proposed the use of the distinct scalings of bipartite fluctuations to detect conventional Landau phase transitions.
@@
Despite decades of study, understanding unconventional quantum phases and phase transitions that are beyond Landau’s symmetry-breaking paradigm remains a central problem in condensed matter physics. In recent years, conceptual breakthroughs [14, 15] have been achieved by employing extended operators to define generalized symmetries and to characterize phases and phase transitions (see Refs. [16, 17] and references therein). For example, abelian topological orders can be understood in terms of the condensation of 1-dimensional objects, leading to the spontaneous breaking of discrete $1$-form symmetries. Similarly, the Coulomb phase of electrodynamics can be interpreted as a spontaneous symmetry-breaking phase of $\textrm{U}(1)$ 11-form symmetry, where gapless photons serve as Goldstone modes. This helps establish a generalized Landau symmetry paradigm. In this case, the criterion for determining whether the phase is symmetric or not is given by the scaling of disorder operators (or Wilson loops under duality). In this context, the concept of bipartite fluctuations is again closely related, which can be identified as the $\textrm{U}(1)$ disorder operator under the small-angle limit [18, 19].
@@
While extended operators play an important role in the conceptual understanding of phases of matter, recent studies have also shed light on intriguing quantitative aspects of disorder operators (and related bipartite fluctuations), particularly their scaling behaviors at conformally invariant quantum critical points [20, 18, 21, 19, 22, 23, 24, 25, 26, 27]. For instance, when considering the configuration of the subsystem as depicted in Fig. 1, bipartite fluctuations exhibit a universal corner contribution with logarithmic scaling. This contribution has a universal angle dependence and is directly proportional to the current central charge $C_{J}$ of the conformal field theory (CFT) [20, 18, 19, 22]. Given the sometimes uncertain fate of proposed lattice realizations for exotic quantum critical points, numerical simulations, such as quantum Monte Carlo simulations, can assist in identifying whether these lattice models correspond to unitary CFTs by examining the sign of the universal corner contribution [23, 24, 25, 26, 27]. Moreover, Ref. [22] showed that the corner term of bipartite fluctuations in certain gapped systems, such as isotropic integer quantum Hall (QH) states and Laughlin states, is determined by the DC Hall conductivity.
@@
Another notable example that has appeared in literature is the bandwidth-tuned transition from a Landau Fermi liquid (FL) to a composite Fermi liquid (CFL) describing the half-filled Landau level. The theoretical possibility of a continuous transition was originally proposed by Ref. [56] and later refined by Ref. [57] targeting moiré materials. The CFL phases at zero magnetic field were predicted to occur in twisted $\textrm{MoTe}_{2}$ at half and three-quarter fillings [58, 59]. On the experimental side, evidence for the CFL phase and the CFL-FL transition has been reported in both twisted $\textrm{MoTe}_{2}$ bilayers [60, 61] and rhombohedral multilayer graphene [62].
A general criterion [63, 64, 65, 66] distinguishing charge insulators from metals is the presence of a finite localization tensor, i.e., finite polarization fluctuations. Accordingly, as will be illustrated below, the CFL may be viewed as a “charge insulator,” and the CFL-FL transition can be interpreted as a “metal-insulator transition” in a broad sense. All of the above-mentioned “metal-insulator transitions” share a common theoretical framework, which we will review in Sec. III, where electrons are fractionalized into fermionic and bosonic partons. The fermion sector forms a stable Fermi-surface state, while the boson sector contains a single relevant operator that drives the transition from a superfluid to a gapped state. Crucial insight into the possible transitions comes from the filling constraints of the bosons under translation symmetry and the Lieb-Schultz-Mattis theorem [55, 57]. Without better terminology, we refer to this type of quantum critical points as critical Fermi surfaces, following Refs. [51, 52].
Figure 1: Real-space subregion $\mathcal{A}$ with a single corner of opening angle $0<\theta<\pi$ and linear size $L$.
@@
In contrast to conformally invariant quantum critical points, many powerful theoretical tools, such as the conformal bootstrap [68, 69], are no longer applicable to strongly correlated metals. However, it is theoretically feasible in Monte Carlo simulations [23, 24, 25, 26, 27] to extract the subleading universal corner term of $\mathcal{F}$ using the method proposed in Ref. [70]. We anticipate that our findings, which establish a connection between the universal data $C_{\rho}$ of the critical points and charge fluctuations which are numerically and experimentally accessible, could aid future studies in identifying the existence and lattice realizations of these exotic quantum critical points in metals.
@@ II Bipartite Fluctuations
In this section, we begin with a brief introduction to the preliminaries of bipartite fluctuations in various quantum states of matter, along with a general criterion for distinguishing charge insulators from metals. We then summarize our main findings on how bipartite fluctuations behave across a class of bandwidth-tuned continuous “metal-insulator transitions.”
@@ II.1 Preliminaries I
For any many-body system with a $\textrm{U}(1)$ global symmetry in $2+1$ dimensions, the bipartite fluctuations associated with a spatial subregion $\mathcal{A}$ are defined by [^2] See App. B for our conventions on the imaginary-time response function $\Pi^{\mu\nu}$ associated with the conserved Noether current $J^{\mu}$.
$\mathcal{F}_{\mathcal{A}}=\int_{\mathcal{A}}\textrm{d}^{2}\bm{x}\int_{\mathcal{A}}\textrm{d}^{2}\bm{y}\>\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x}-\bm{y}),$ (1)
where $\Pi^{\tau\tau}$ denotes the equal-time density-density correlation function, i.e., the static structure factor.
In this introductory section, we focus on isotropic systems for simplicity and review several known results. More general and anisotropic systems have been discussed in Ref. [20] and later in Refs. [71, 72], where certain non-universal aspects are explored.
In App. D, we evaluate $\mathcal{F}_{\mathcal{A}}$ assuming a generic power-law instantaneous charge correlation in two dimensions
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})=\frac{-C_{0}}{|\bm{x}|^{\alpha}}.$ (2)
The resulting scaling behaviors are summarized in Tab. 1.
@@
$2<\alpha<3$ $L^{4-\alpha}$ $L^{4-\alpha}$
@@
$3<\alpha<4$ $L/\epsilon^{\alpha-3}$ $L^{4-\alpha}$
@@
$\alpha>4$ $L/\epsilon^{\alpha-3}$ 0
Table 1: Scaling behaviors of the bipartite fluctuations $\mathcal{F}_{\mathcal{A}}$, based on the instantaneous density-density correlation Eq. (2). Here, $L$ denotes the linear size of the subregion $\mathcal{A}$, and $\epsilon$ is a gauge-invariant UV cutoff introduced in the vortex theory (see App. D). Notable features include: the leading-order term follows boundary-law scaling for $\alpha>3$; a universal subleading logarithmic term appears only when $\alpha=4$ and the geometry of $\mathcal{A}$ includes sharp corners; and for $\alpha>4$, the $\epsilon$-independent subleading term vanishes in the large-$L$ limit.
The simplest compressible state is the free Fermi gas. Using the free-fermion propagator, one can directly compute $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})$ in real space and the static structure factor $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ in momentum space
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})$ $=\frac{k_{F}}{4\pi^{3}}\frac{-1}{|\bm{x}|^{3}},$
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ $=\frac{k_{F}|\bm{k}|}{2\pi^{2}},$ (3)
where $k_{F}$ denotes the Fermi momentum. According to Tab. 1, we observe the scaling $\mathcal{F}\sim L\log(L)$. This is a well-known result closely related to the entanglement entropy of free fermions [73, 74, 75, 76, 9, 8, 4, 77]. Recent studies [78, 79] have further explored related calculations addressing the role of Fermi-surface topology in free-fermion systems. In App. A, we reproduce Eq. (3) based the $\textrm{LU}(1)$ anomaly of Fermi surface states [46], which remains valid even in the presence of strong forward scattering that preserves $\textrm{LU}(1)$ symmetry. This approach also provides a geometric interpretation [80] of the static structure factor in terms of the area difference induced by shifting the Fermi surface, as illustrated in Fig. 2.
Figure 2: Geometric interpretation of the static structure factor $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{q})$ for Fermi-surface states with LU(1) symmetry, where the shaded area indicates its value.
Another simple compressible state is provided by the superfluid phase. Let $\phi$ denote the $\mathrm{U}(1)$ order parameter. The gapless Goldstone mode $\theta$ describes the phase fluctuations of $\phi=\rho_{s}e^{\mathtt{i}\theta}$, where $\rho_{s}$ is the superfluid stiffness. Power counting of the action $\int\textrm{d}^{3}x(\partial\theta)^{2}$ yields a scaling dimension of $\Delta[\partial\theta]=3/2$ for the $\textrm{U}(1)$ charge density, implying that the equal-time density-density correlation decays as $|\bm{x}|^{-3}$. Once again, one finds the scaling of bipartite fluctuations $\mathcal{F}\sim L\log(L)$.
@@
$\mathcal{F}=\#L-\frac{C_{J}}{2}\textrm{f}(\theta)\log(L)+\ldots.$ (5)
where #\# is a non-universal number depending on the UV cut-off, and the universal angle dependence
@@
applies to any 2+1D CFTs. This corner contribution has also been previously discussed in the context of non-interacting Dirac systems [20]. However, the question remains whether the universal behavior described in Eq. (5) could also exist at certain quantum critical points lacking conformal symmetry.
Another standard class of incompressible states is insulators with a charge gap. In these systems, the equal-time density-density correlation decays exponentially, leading to a boundary-law scaling of bipartite fluctuations. An intriguing observation from Ref. [22] is that certain gapped QH insulators with continuous translational and rotational symmetries exhibit a universal corner contribution when the subregion geometry includes a sharp corner, as in Fig. 1:
$\mathcal{F}=\#L-\frac{|\sigma_{xy}|}{2\pi}\textrm{f}(\theta)+\ldots.$ (7)
The angle dependence is once again given by Eq. (6), and $\sigma_{xy}=\nu/(2\pi)$ represents the DC Hall conductivity, where $\nu$ denotes the filling factor. For integer fillings $\nu\in\mathbb{Z}$, the corner term can be derived analytically, while for the fractional QH insulator at $\nu=1/3$, it has been verified using Monte Carlo simulations based on the Laughlin wavefunction [22]. It would be interesting to explore whether a similar corner term exists in compressible QH states at even-denominator fillings and to understand the conditions under which the formula in Eq. (7) holds.